OpenStudy (anonymous):

A company discovers that to produce 520 new electronic parts, it will cost $42960. To produce 820 new electronic parts, it will cost $64860. Compute the slope of the costs and choose the most accurate statement from the following: Costs are increasing at a rate of $75 per item. Costs are decreasing at a rate of $73 per item. Costs are decreasing at a rate of $75 per item. Costs are decreasing at a rate of $70 per item. Costs are increasing at a rate of $70 per item. Costs are increasing at a rate of $73 per item.

OpenStudy (anonymous):

They give you two points (520, 42960) and (820,64860) Do you know how to compute the slope from here?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Then there you go.

OpenStudy (anonymous):

But then the answer is 52, which isn't a choice.

OpenStudy (anonymous):

I get 73 how are you computing the slope?

OpenStudy (anonymous):

y2-y1/x2-x1

OpenStudy (anonymous):

Yep, that's correct. Try again it is just an arithmetic error

OpenStudy (anonymous):

64860-49260 = 15600 & 820 - 520 =300 15600/300 = 52?